# C. Röver, Q. Wu, A. Loos, T. Friede

### v1

@@ Abstract
[Summary] To appreciate intervention effects on rare events, meta-analysis techniques are commonly applied in order to assess the accumulated evidence. When it comes to adverse effects in clinical trials, these are often most adequately handled using survival methods. A common-effect model that is able to process data in commonly quoted formats in terms of hazard ratios has been proposed for this purpose by Holzhauer (Stat. Med. 2017; 36(5):723–737). In order to accommodate potential heterogeneity between studies, we have extended the model by Holzhauer to a random-effects approach. The Bayesian model is described in detail, and applications to realistic data sets are discussed along with sensitivity analyses and Monte Carlo simulations to support the conclusions.
@@ 2.2.2 Oncology data
[…]
@@ 3.1 The common-effect model by Holzhauer
For the log-HR $\varphi$, we propose to employ either a $\mathrm{Cauchy}(0.0,2.5)$ distribution as weakly informative prior31 or a $\mathrm{Cauchy}(0.0,0.37)$, which puts 90% of the prior probability mass within $[-\log(10),\log(10)]$ reflecting the belief that extremely large drug effects with odds ratios above 10 or below 0.1 are rare. 32, 33
@@ 3.2 The proposed random-effects model
For the overall log-HR $\varphi$, it may make sense to adopt the same priors ($\mathrm{Cauchy}(0.0,2.5)$ or $\mathrm{Cauchy}(0.0,0.37)$) as in Section 3.1. A $\mathrm{Halfnormal}(100)$ distribution may be used as a vague prior on the scale $\eta$, essentially as a conservative specification mimicking an (improper) uniform prior.27 Otherwise, a $\mathrm{Halfnormal}(0.5)$ distribution is used as a weakly informative prior on $\eta$, especially when there is a small number of trials in the main meta-analysis.36
@@ 3.3 Implementation of the models
The Bayesian analyses are carried out using Markov chain Monte Carlo (MCMC) methods.37 The CE model was originally implemented using the the proc MCMC procedure within the SAS software.3 The random-effects extension was now implemented in R (version 4.3.2) and Stan,38 as well as the rstan39 (version 2.32.5) and RBesT30(version 1.7-3) packages. Derivation of MAP priors was implemented using the mclust, teigen and RBesT functions.12
@@ 4.1 Rosiglitazone data
For the purpose of comparing common-effect (CE) and random-effects (RE) models, we apply both on the rosiglitazone data set and visualize the point estimate for the log-HR ($\varphi$) along with 95% equal-tailed credible intervals (CIs) in Figure 3. One can see that RE models yield wider CIs than CE models in all pairs of comparisons. Because RE models consider between-trial heterogeneity in addition, the estimation uncertainty is larger than in CE models. Another finding is that point estimates of the HR in RE models in this case are generally smaller than in CE models. Some variation in point estimates is expected, because the weighting of different trials in a meta-analysis will change when moving from a CE model to a RE model.40 CE models yielding larger point estimates mean that trials with larger HR tend to get greater weights, whereas when weights become more even in a RE model, the impact of these trials is reduced, and the point estimate, as a weighted average of all trials’ HRs, becomes smaller. When checking the rosiglitazone data, we indeed find that trials with large (positive) effect mostly coincide with large sample size; so this tendency has a reasonable explanation.
@@ 8 $\mathrm{Halfnormal}(100)$
Figure 4: Sensitivity analysis for RE models with varying priors on $\eta$ in case of a large number of trials in the rosiglitazone example. Blue lines correspond to models without borrowing from historical trials; red lines correspond to models with borrowing from historical trials. Each line represents the 95% equal-tailed CI for $\varphi$, the point estimate is indicated by a box on the line.
@@
Figure 5: Sensitivity analyses for RE models with varying priors on the heterogeneity $\eta$ in case of a small number of trials in the rosiglitazone example. Red and blue lines correspond to models with and without borrowing from historical trials. Each line represents the 95% equal-tailed CI for $\varphi$, the point estimate is indicated by a square.
@@ 4.2 Oncology data
Figure 7: Sensitivity analysis for RE models with varying priors for the heterogeneity $\eta$ in the oncology example. Each line represents the 95% equal-tailed CI for the overall HR $\varphi$, the point estimate is indicated by a square.
@@ 5.1 Aims
To investigate the operating characteristics of RE models in comparison to the CE model, we performed two simulation studies motivated by the two data examples introduced in Section 2.2. We will in particular consider coverage probabilities, type I error and power of the different methods in a limited range of scenarios.
@@ 5.2.1 Setup
Data are generated based on the model described in Section 3.2. We assume the historical trials share the same event rate ($\lambda_{i0}=0.5$) and drop-out rate ($\mu_{i0}=\mu_{i1}=0.5$) with the main meta-analysis. Also, the proportions of fatal events were constant ($q_{0}=q_{1}=0.35$). We simulate the IPD using parameters shown in Table 3, and then sum over the patients in each trial-arm to derive the corresponding aggregate data. The major difference between scenarios is in their underlying hazard ratio (HR). We assume either null or positive treatment effects. Scenarios 2–4 and 6–8 assume random treatment effects, namely, the trial-specific log-HR $\varphi_{i}$ following a $\mathrm{Normal}(\varphi,\eta^{2})$ distribution (18). When determining the amount of between-trial heterogeneity, we first consider the mean of a $\mathrm{Halfnormal}(0.5)$ distribution ($\operatorname{E}[\eta]=0.40$), as this is considered as a reasonable prior for $\eta$ when the number of trials is small.36 Then the value of $\eta$ is further varied to be half or twice as much, to capture situations with small and large between-trial heterogeneity. In addition, scenarios 1 and 5 assume no between-trial heterogeneity ($\eta=0$, homogeneity, CE assumption). For each scenario, we perform 1000 simulations.
Table 3: Parameter settings used in the simulations motivated by the rosiglitazone application. In each scenario, $I=6$ studies and in addition $H=12$ historical (control) studies were used. Scenarios 2–4 and 6–8 correspond to a RE setting ($\eta>0$), while scenarios 1 and 5 implement a CE assumption ($\eta=0$). For more details, see Section 5.2.1.
scenario log-HR $\varphi$ heterogeneity $\eta$
1 0.00 0.00 (none)
2 0.00 0.20 (small)
3 0.00 0.40 (moderate)
4 0.00 0.80 (large)
5 0.25 0.00 (none)
6 0.25 0.20 (small)
7 0.25 0.40 (moderate)
8 0.25 0.80 (large)
@@ 5.2.2 Results
Table 4 shows the coverage rate of 95% credible intervals (CIs) for the overall HR $\varphi$ for all four models in each scenario. If we compare the CE models against the RE models, we find that in all scenarios CE models have lower CI coverage rates than RE models, because RE models account for between-trial heterogeneity in addition to sampling error, leading to wider CIs and increased coverage. For scenarios 2–4 and 6–8, RE models should perform better, as they should be able to account for REs implemented in the data generation. This is supported by the observation from Table 4; for scenarios 1 and 5, since the data generating mechanism assumes no heterogeneity between trials, the CE model should perform better. Consistently, we indeed find the coverage rate for common effect model in scenarios 1 and 5 to be very close to 95 %, whereas the CI coverage rates of RE models are greater, suggesting that RE models are more conservative in these cases.
@@
scenario prior common effect (CE) random effects (RE)
1 vague 94.5 98.8
@@
If we further compare the models with and without borrowing from historical trials (using vague priors or MAP priors), we find that in scenarios 2–4 and 6–8, vague models provide higher CI coverage rate than MAP models, while scenarios 1 and 5 show vague models with lower coverage than MAP models. Such trends can be attributed to the difference in point estimates of $\varphi$ between vague and MAP models. A further investigation shows that for scenarios 2–4 and 6–8, the point estimates of vague and MAP models are very close to each other. Since MAP models have narrower CIs due to additional information borrowed from historical trials, vague models have lower CI coverage than MAP models. However, for scenarios 1 and 5, there is a clear discrepancy between the point estimates of MAP and vague models. The point estimates of MAP models are closer to the true value. Therefore, although MAP models have narrower CIs, they still have higher coverage due to the more precise point estimates.
Apart from CI coverage, we further investigate the type I error rate for scenarios under the null hypothesis, and power for scenarios under alternative hypotheses, respectively. Table 5 shows the type I error rates of scenarios 1–4 which assume $\varphi$ to be zero (i.e., no adverse effect). For the CE scenario (scenario 1), CE models control the type I error rate at around $5\%$, but RE models result in conservative type I error rates ($>5\%$). For the other scenarios, in contrast to CE models, RE models can better control the type I error rate, because random effects models tend to have wider CIs and thus lower rejection rates.
@@
model scenario: 1 2 3 4 5 6 7 8
CE-vague 05.5 10.2 27.8 54.1 51.4 55.4 61.5 73.0
CE-MAP 03.3 12.7 31.6 56.2 79.2 73.0 68.6 74.8
RE-vague 01.2 01.7 02.9 02.8 29.0 22.6 12.4 07.0
RE-MAP 00.7 02.4 05.2 07.8 47.5 35.7 22.0 16.6
Table 5 also shows the power for scenarios 5–8 which assume $\varphi$ to be different from zero. We focus on the models successfully controlling the type I error rate in the same table. For the RE models in Table 5, MAP models have greater power than vague models, because MAP models with additional historical information yield narrower CIs and thus higher rejection rates. Since the underlying data generating mechanism assumes a CE in scenario 5, CE models outperform RE models in terms of power, as expected.
To summarize, in the presence of between-trial heterogeneity, RE models tend to perform better than CE models in terms of CI coverage and type I error control. In case of homogeneity, RE models are too conservative, leading to too high CI coverage, too low type I error rate and low power. The situation could be mitigated by choosing a smaller scale for the prior of $\eta$ in such cases. For example, for scenarios 1 and 5, if we set the prior of $\eta$ to be $\mathrm{Halfnormal}(0.25)$ instead of $\mathrm{Halfnormal}(100)$ for vague models and $\mathrm{Halfnormal}(0.5)$ for MAP models, the type I error rate of RE models for scenario 1 increases from around 1% to around 2%, while the power of RE models for scenario 5 increases from 29.0% to 38.5% for the vague model and from 47.5% to 58.4% for the MAP model, respectively.
@@ 5.3.1 Setup
Since oncology trials have many unique features, we considered it worthwhile to separately conduct a simulation study to investigate the behavior of the proposed models under such circumstances. The simulation design is motivated by the oncology data example introduced in Section 2.2.2. We fit the RE model using vague priors to the data and designed simulation scenarios listed in Table 6 based on the derived estimates. The major difference among scenarios still lies in the treatment effects. Following the same structure as the rosiglitazone simulation, we make the treatment effects be random for most scenarios and fixed for the scenarios 1 and 5. Either a null or positive treatment effect is assumed for the value of the HR $\varphi$. Between-trial heterogeneity varies from 0.0 (homogeneity), over 0.2 (small) and 0.4 (moderate) to 0.8 (large).
Table 6: Parameter settings used in the simulations motivated by the oncology application. In each scenario, $I=9$ studies are considered, but no historical studies ($H=0$). Scenarios 2-4 and 6–8 correspond to a RE setting ($\eta>0$), while scenarios 1 and 5 implement a CE assumption ($\eta=0$). For more details, see Section 5.3.1.
scenario log-HR $\varphi$ heterogeneity $\eta$
1 0.00 0.00 (none)
2 0.00 0.20 (small)
3 0.00 0.40 (moderate)
4 0.00 0.80 (large)
5 0.50 0.00 (none)
6 0.50 0.20 (small)
7 0.50 0.40 (moderate)
8 0.50 0.80 (large)
@@ 5.3.2 Results
Table 7 shows the coverage rate of 95% equal-tailed CIs of both CE and RE models in each scenario. For all scenarios, CE models have lower CI coverage than RE models, as expected, since RE models additionally consider between-trial heterogeneity and thus have wider CIs. For RE scenarios (scenarios 2–4 and 6–8), RE models should give CI coverage closer to 95% than CE models. But for CE scenarios (scenarios 1 and 5), the CI coverage of CE models should be closer to 95% than RE models. Results in Table 7 match our expectation well. Focusing on scenarios 1–4 in Table 7, we observe a decrease in CI coverage as the heterogeneity $\eta$ increases, for both CE and RE models. The same trend can be observed for scenarios 4–8. The obvious explanation is that when the true value of $\eta$ is larger, CE models tend to underestimate the total variation more, thus the resulting CI is much narrower, and coverage decreases. For RE models, scenarios 3 and 7 have CI coverage rate quite close to 95% because the prior for $\eta$ ($\mathrm{Halfnormal}(0.5)$) is consistent with its true value ($\eta=0.4$). For scenarios 2 and 6, the prior expectation of $\eta$ is larger than its true value, thus the between-trial heterogeneity is overestimated, resulting in too wide CIs and an increased (conservative) coverage rate. For scenarios 4 and 8, the prior expectation on $\eta$ is lower than its true value, so the between-trial heterogeneity gets underestimated, leading to too narrow CIs and thus too low coverage rate.
@@
scenario prior common effect (CE) random effects (RE)
1 vague 95.9 98.5
2 vague 86.1 96.8
3 vague 68.8 94.3
4 vague 40.7 89.9
5 vague 94.7 98.4
6 vague 88.6 98.3
7 vague 71.5 95.6
8 vague 43.6 91.1
As in the previous simulation (Section 5.2), we also investigate other statistical properties such as the type I error and power. Table 8 shows the type I error rates of no-treatment-effect scenarios 1–4. For scenario 1, which assumes no between-trial heterogeneity, the CE model can control the type I error rate at below 5%, while the RE model results in too conservative type I error rate. For all four scenarios, RE models have lower type I error rates than CE models. We focus on cases where type I error is controlled in Table 8 and consider their power. Under a CE scenario (scenario 5), the CE model achieves high power (above 90%), while the RE model attains lower power due to their larger CI width. For RE models, as between-trial heterogeneity increases from scenario 5 to scenario 6, 7 and 8, the power decreases. This is because higher heterogeneity leads to lower signal-to-noise ratio. As a result, the signal is more difficult to identify, and hence correct rejection is more difficult.
@@
model scenario: 1 2 3 4 5 6 7 8
CE-vague 4.1 13.9 31.2 59.3 91.1 86.0 74.1 60.7
RE-vague 1.5 03.2 05.7 10.1 79.0 72.5 57.5 34.8
In summary, analogous trends are observed in the oncology simulation as in the rosiglitazone simulation. The only difference we find is in the absolute magnitude of power. Comparing Tables 5 and 8, the power for the oncology simulation is much higher than in the rosiglitazone simulation. This may most likely be attributed to the larger effect ($\varphi$) in the oncology simulations, leading to easier rejection and thus higher power in the oncology simulation.

### v2

@@ Abstract
[Summary] To investigate intervention effects on rare events, meta-analysis techniques are commonly applied in order to assess the accumulated evidence. When it comes to adverse effects in clinical trials, these are often most adequately handled using survival methods. A common-effect model that is able to process data in commonly quoted formats in terms of hazard ratios has previously been proposed for this purpose. In order to accommodate potential heterogeneity between studies, we have extended the model by Holzhauer to a random-effects approach. The Bayesian model is described in detail, and applications to realistic data sets are discussed along with sensitivity analyses and Monte Carlo simulations to support the conclusions.
@@ 2.2.2 Oncology data
Figure 2: Typical scenarios for adverse event follow‐up in clinical trials (adopted from Figure 1 in Unkel et al.(2019)2). The abbreviations are: AE, adverse event; TEAE, treatment-emergent adverse event; FU, follow-up; Saf-FU, safety follow-up; EoT, end of treatment; V0, visit at trial onset; V11,…,Vnn, visits during treatment. First occurrences of AEs are marked by triangles.
@@ 3.1 The common-effect model by Holzhauer
For the log-HR $\varphi$, we propose to employ either a $\mathrm{Cauchy}(0.0,2.5)$ distribution as weakly informative prior31 or a $\mathrm{Cauchy}(0.0,0.37)$, which puts 90% of the prior probability mass within $\log(10),\log(10)]$ reflecting the belief that extremely large drug effects with odds ratios above 10 or below 0.1 are rare. 32, 33
@@ 3.2 The proposed random-effects model
For the overall log-HR $\varphi$, it may make sense to adopt the same priors ($\mathrm{Cauchy}(0.0,2.5)$ or $\mathrm{Cauchy}(0.0,0.37)$) as in Section 3.1. A $\mathrm{Halfnormal}(100)$ distribution may be used as a vague prior on the scale $\eta$, essentially as a conservative specification mimicking an (improper) uniform prior.27 Otherwise, a $\mathrm{Halfnormal}(0.5)$ distribution is used as a weakly informative prior on $\eta$, especially when there is a small number of trials in the main meta-analysis.28
@@ 3.3 Implementation of the models
The Bayesian analyses are carried out using Markov chain Monte Carlo (MCMC) methods.36 The CE model was originally implemented using the the proc MCMC procedure within the SAS software, while a Stan implementation has also been used for simulations.3, 37 The random-effects extension was now implemented in R (version 4.3.2) and Stan,38 as well as the rstan39 (version 2.32.5) and RBesT30(version 1.7-3) packages. Derivation of MAP priors was implemented using the mclust, teigen and RBesT functions.12 Stan implements a Hamiltonian Monte Carlo sampler; each time, $5$ chains were run in parallel, to ensure proper burn-in, the initial 10 00010\,000 samples of each chain were discarded, and to reduce autocorrelations, only every $5$th sample of each chain was kept. Simulations eventually lasted approximately 10 00010\,000 samples for each iteration, and convergence was checked using the rank-based $\hat{R}$.40 Prior sensitivity is quantified using the $D_{\mathrm{CJS}}$ metric as implemented in the priorsense R package, where smaller values indicate less sensitivity of results to changes in the corresponding parameter’s prior specification.41, 42
@@ 4.1 Rosiglitazone data
For the purpose of comparing common-effect (CE) and random-effects (RE) models, we apply both on the rosiglitazone data set and visualize the point estimate for the log-HR ($\varphi$) along with 95% equal-tailed credible intervals (CIs) in Figure 3. One can see that RE models yield wider CIs than CE models in all pairs of comparisons. Because RE models consider between-trial heterogeneity in addition, the estimation uncertainty is larger than in CE models. Another finding is that point estimates of the HR in RE models in this case are generally smaller than in CE models. Some variation in point estimates is expected, because the weighting of different trials in a meta-analysis will change when moving from a CE model to a RE model.43 CE models yielding larger point estimates mean that trials with larger HR tend to get greater weights, whereas when weights become more even in a RE model, the impact of these trials is reduced, and the point estimate, as a weighted average of all trials’ HRs, becomes smaller. When checking the rosiglitazone data, we indeed find that trials with large (positive) effect mostly coincide with large sample size; so this tendency has a reasonable explanation.
@@ 8 $\mathrm{Halfnormal}(100)$
Figure 4: Sensitivity analysis for RE models with varying priors on $\eta$ in case of a large number of trials in the rosiglitazone example. Blue lines correspond to models without borrowing from historical trials; red lines correspond to models with borrowing from historical trials. Each line represents the 95% equal-tailed CI for $\varphi$, the point estimate is indicated by a box on the line. Prior sensitivity (with respect to $\eta$) is quantified in terms of the $D_{\mathrm{CJS}}$ metric.
@@
Figure 5: Sensitivity analyses for RE models with varying priors on the heterogeneity $\eta$ in case of a small number of trials in the rosiglitazone example. Red and blue lines correspond to models with and without borrowing from historical trials. Each line represents the 95% equal-tailed CI for $\varphi$, the point estimate is indicated by a square. Prior sensitivity (with respect to $\eta$) is quantified in terms of the $D_{\mathrm{CJS}}$ metric.
In both scenarios (Figures 4 and 5), prior sensitivity with respect to the heterogeneity $\eta$ was investigated in terms of the $D_{\mathrm{CJS}}$ metric.41 Sensitivity tends to be smallest for a conventional choice of a prior scale of $0.5$, suggesting that priors favouring more extreme (small or large) heterogeneities may imply a prior/data conflict.
@@ 4.2 Oncology data
Figure 7: Sensitivity analysis for RE models with varying priors for the heterogeneity $\eta$ in the oncology example. Each line represents the 95% equal-tailed CI for the overall HR $\varphi$, the point estimate is indicated by a square. Prior sensitivity (with respect to $\eta$) is quantified in terms of the $D_{\mathrm{CJS}}$ metric.
Again, the $\eta$ prior sensitivity (as quantified through the $D_{\mathrm{CJS}}$ metric) tends to be smallest for a weakly informative $\mathrm{Halfnormal}(0.5)$ prior.
@@ 5.1 Aims
To investigate the operating characteristics of RE models in comparison to the CE model, we performed two simulation studies motivated by the two data examples introduced in Section 2.2. We will in particular consider coverage probabilities, type I error and power of the different methods in a limited range of scenarios. Additional simulation results investigating departures from the exponential distribution assumption (1) are provided in Appendix C.
@@ 5.2.1 Setup
Data are generated based on the model described in Section 3.2. We assume the historical trials share the same event rate ($\lambda_{i0}=0.5$) and drop-out rate ($\mu_{i0}=\mu_{i1}=0.5$) with the main meta-analysis. Also, the proportions of fatal events were constant ($q_{0}=q_{1}=0.35$). We simulate the IPD using parameters shown in Table 3, and then sum over the patients in each trial-arm to derive the corresponding aggregate data. The major difference between scenarios is in their underlying hazard ratio (HR). We assume either null or positive treatment effects. Scenarios 2–4 and 6–8 assume random treatment effects, namely, the trial-specific log-HR $\varphi_{i}$ following a $\mathrm{Normal}(\varphi,\eta^{2})$ distribution (18). When determining the amount of between-trial heterogeneity, we first consider the mean of a $\mathrm{Halfnormal}(0.5)$ distribution ($\operatorname{E}[\eta]=0.40$), as this is considered as a reasonable prior for $\eta$ when the number of trials is small.28 Then the value of $\eta$ is further varied to be half or twice as much, to capture situations with small and large between-trial heterogeneity. In addition, scenarios 1 and 5 assume no between-trial heterogeneity ($\eta=0$, homogeneity, CE assumption). For each scenario, we perform 1000 simulations.
@@ 5.2.2 Results
Table 3 shows the coverage rate of 95% credible intervals (CIs) for the overall HR $\varphi$ for all four models in each scenario. If we compare the CE models against the RE models, we find that in all scenarios CE models have lower CI coverage rates than RE models, because RE models account for between-trial heterogeneity in addition to sampling error, leading to wider CIs and increased coverage. For scenarios 2–5 and 7–10, RE models should perform better, as they should be able to account for REs implemented in the data generation. This is supported by the observation from Table 3; for scenarios 1 and 6, since the data generating mechanism assumes no heterogeneity between trials, the CE model should perform better. Consistently, we indeed find the coverage rate for common effect model in scenarios 1 and 6 to be very close to 95 %, whereas the CI coverage rates of RE models are greater, suggesting that RE models are more conservative in these cases. The RE models’ coverage then decreases with increasing heterogeneity, and (as expected) may tend towards undercoverage if the heterogeneity setting approaches the upper limit of the a priori expected range (scenarios 5 and 10).
@@
simulation settings coverage probability (%)
scenario log-HR $\varphi$ heterogeneity $\eta$ prior common effect (CE) random effects (RE)
1 0.00 0.00 (none) vague 94.5 98.8
@@
If we further compare the models with and without borrowing from historical trials (using vague priors or MAP priors), we find that in scenarios 2–5 and 7–10, vague models provide higher CI coverage rate than MAP models, while scenarios 1 and 6 show vague models with lower coverage than MAP models. Such trends can be attributed to the difference in point estimates of $\varphi$ between vague and MAP models. A further investigation shows that for scenarios 2–5 and 7–10, the point estimates of vague and MAP models are very close to each other. Since MAP models have narrower CIs due to additional information borrowed from historical trials, vague models have lower CI coverage than MAP models. However, for scenarios 1 and 6, there is a clear discrepancy between the point estimates of MAP and vague models. The point estimates of MAP models are closer to the true value. Therefore, although MAP models have narrower CIs, they still have higher coverage due to the more precise point estimates.
Apart from CI coverage, we further investigate the type I error rate for scenarios under the null hypothesis, and power for scenarios under alternative hypotheses, respectively. Table 4 shows the type I error rates of scenarios 1–5 which assume $\varphi$ to be zero (i.e., no adverse effect). For the CE scenario (scenario 1), CE models control the type I error rate at around $5\%$, but RE models result in conservative type I error rates ($5\%$). For the other scenarios, in contrast to CE models, RE models can better control the type I error rate, because random effects models tend to have wider CIs and thus lower rejection rates.
@@
model scenario: 1 2 3 4 5 6 7 8 9 10
CE-vague 05.5 7.5 10.2 27.8 54.1 51.4 51.8 55.4 61.5 73.0
CE-MAP 03.3 6.0 12.7 31.6 56.2 79.2 75.6 73.0 68.6 74.8
RE-vague 01.2 1.1 01.7 02.9 02.8 29.0 24.5 22.6 12.4 07.0
RE-MAP 00.7 0.9 02.4 05.2 07.8 47.5 45.2 35.7 22.0 16.6
Table 4 also shows the power for scenarios 6–10 which assume $\varphi$ to be different from zero. We focus on the models successfully controlling the type I error rate in the same table. For the RE models in Table 4, MAP models have greater power than vague models, because MAP models with additional historical information yield narrower CIs and thus higher rejection rates. Since the underlying data generating mechanism assumes a CE in scenario 6, CE models outperform RE models in terms of power, as expected. The accuracy of point estimates for the log-HR is summarized in Table 7 (Appendix B).
To summarize, in the presence of between-trial heterogeneity, RE models tend to perform better than CE models in terms of CI coverage and type I error control. In case of homogeneity, RE models are too conservative, leading to too high CI coverage, too low type I error rate and low power. The situation could be mitigated by choosing a smaller scale for the prior of $\eta$ in such cases. For example, for scenarios 1 and 6, if we set the prior of $\eta$ to be $\mathrm{Halfnormal}(0.25)$ instead of $\mathrm{Halfnormal}(100)$ for vague models and $\mathrm{Halfnormal}(0.5)$ for MAP models, the type I error rate of RE models for scenario 1 increases from around 1% to around 2%, while the power of RE models for scenario 5 increases from 29.0% to 38.5% for the vague model and from 47.5% to 58.4% for the MAP model, respectively.
@@ 5.3.1 Setup
Since oncology trials have many unique features, we considered it worthwhile to separately conduct a simulation study to investigate the behavior of the proposed models under such circumstances. The simulation design is motivated by the oncology data example introduced in Section 2.2.2. We fit the RE model using vague priors to the data and designed simulation scenarios listed in Table 5 based on the derived estimates. The major difference among scenarios still lies in the treatment effects. Following the same structure as the rosiglitazone simulation, we make the treatment effects be random for most scenarios and fixed for the scenarios 1 and 5. Either a null or positive treatment effect is assumed for the value of the HR $\varphi$. Between-trial heterogeneity varies from 0.0 (homogeneity), over 0.1 (small), 0.2 (moderate) and 0.4 (substantial) to 0.8 (large).
@@ 5.3.2 Results
Table 5 shows the coverage rate of 95% equal-tailed CIs of both CE and RE models in each scenario. For all scenarios, CE models have lower CI coverage than RE models, as expected, since RE models additionally consider between-trial heterogeneity and thus have wider CIs. For RE scenarios (scenarios 2–5 and 7–10), RE models should give CI coverage closer to 95% than CE models. But for CE scenarios (scenarios 1 and 6), the CI coverage of CE models should be closer to 95% than RE models. Results in Table 5 match our expectation well. Focusing on scenarios 1–5 in Table 5, we observe a decrease in CI coverage as the heterogeneity $\eta$ increases, for both CE and RE models. The same trend can be observed for scenarios 6–10. The obvious explanation is that when the true value of $\eta$ is larger, CE models tend to underestimate the total variation more, thus the resulting CI is much narrower, and coverage decreases. For RE models, scenarios 4 and 8 have CI coverage rate quite close to 95% because the prior for $\eta$ ($\mathrm{Halfnormal}(0.5)$) is consistent with its true value ($\eta=0.4$). For scenarios 3 and 8, the prior expectation of $\eta$ is larger than its true value, thus the between-trial heterogeneity is overestimated, resulting in too wide CIs and an increased (conservative) coverage rate. For scenarios 5 and 10, the prior expectation on $\eta$ is lower than its true value, so the between-trial heterogeneity gets underestimated, leading to too narrow CIs and thus too low coverage rate.
@@
simulation settings coverage probability (%)
scenario log-HR $\varphi$ heterogeneity $\eta$ prior common effect (CE) random effects (RE)
1 0.00 0.00 (none) vague 95.9 98.5
2 0.00 0.10 (small) vague 93.2 98.0
3 0.00 0.20 (moderate) vague 86.1 96.8
4 0.00 0.40 (substantial) vague 68.8 94.3
5 0.00 0.80 (large) vague 40.7 89.9
6 0.50 0.00 (none) vague 94.7 98.4
7 0.50 0.10 (small) vague 92.7 97.7
8 0.50 0.20 (moderate) vague 88.6 98.3
9 0.50 0.40 (substantial) vague 71.5 95.6
10 0.50 0.80 (large) vague 43.6 91.1
As in the previous simulation (Section 5.2), we also investigate other statistical properties such as the type I error and power. Table 6 shows the type I error rates of no-treatment-effect scenarios 1–5. For scenario 1, which assumes no between-trial heterogeneity, the CE model can control the type I error rate at below 5%, while the RE model results in too conservative type I error rate. For all five scenarios, RE models have lower type I error rates than CE models. We focus on cases where type I error is controlled in Table 6 and consider their power. Under a CE scenario (scenario 6), the CE model achieves high power (above 90%), while the RE model attains lower power due to their larger CI width. For RE models, as between-trial heterogeneity increases from scenario 6 to scenario 7, 8, 9 and 10, the power decreases. This is because higher heterogeneity leads to lower signal-to-noise ratio. As a result, the signal is more difficult to identify, and hence correct rejection is more difficult. The accuracy of point estimates for the log-HR is summarized in Table 8 (Appendix B).
@@
model scenario: 1 2 3 4 5 6 7 8 9 10
CE-vague 4.1 6.8 13.9 31.2 59.3 91.1 85.8 86.0 74.1 60.7
RE-vague 1.5 2.0 03.2 05.7 10.1 79.0 75.1 72.5 57.5 34.8
In summary, analogous trends are observed in the oncology simulation as in the rosiglitazone simulation. The only difference we find is in the absolute magnitude of power. Comparing Tables 4 and 6, the power for the oncology simulation is much higher than in the rosiglitazone simulation. This may most likely be attributed to the larger effect ($\varphi$) in the oncology simulations, leading to easier rejection and thus higher power in the oncology simulation.